u(x,t) is solution to heat equation, ,with following parameters for numerical approximation:

0 < x < 2,    0 < t < 0.1, n = 20, m = 100, c =1. Boundary conditions: u(0,t) =0, and u(2,0) = 0.
Initial conditions: u (x,0) =30o for 0<x<=1
                                            0o for 1<x<2
Set the approximate difference equation for this equation.

Do you think this equation converges to a numerical solution.

Continuing with problem 1, calculate u(0.1,0.001) by iteration

Continuing with problem 1, calculate u(0.2,0.001) by iteration
Answer #1

Forward time central difference scheme: 4, +1,5 – 24,, +U:-1,5 – 4,;+1 –Uj (14 ,1,–21., +4,-43] =1:38 U Here, h=2-0 = 0.1 20